Find the hypotenuse of a right triangle with sides 5 inches and 10 inches.
Answer by checkley71(8403)

5^2+10^2=H^2
25=100=H^2
150=H^2
H=SQRT150
H=12.25 INCHES IS THE LENGTH OF THE HYPOTENUSE.